OpenStudy (anonymous):

The expression 4n – 1 is a rule to describe a certain sequence. What are the first three terms of the sequence?

OpenStudy (shamil98):

n = 1 n = 2 n = 3 input those into 4n-1. and evaluate.

OpenStudy (happinessbreaksbones):

beat me to it, shamil.

OpenStudy (anonymous):

For the first term, substitute n = 1. For the second term, substitute n = 2. For the third term, substitute n = 3. See how easy it can be?

OpenStudy (anonymous):

What does that mean?

OpenStudy (anonymous):

Plug in: If n=1 ,so 4(1)-1=3 Then plug in n=2 n=3

OpenStudy (shamil98):

4(1) - 1 = 1st term 4(2) - 1 = 2nd term 4(3) - 1 = 3d term

OpenStudy (anonymous):

It is multiple choice :)

OpenStudy (anonymous):

3, 7, 11 4, 8, 12 3, 6, 9 4, 7, 11

OpenStudy (anonymous):

4(1) - 1 = 1st term 4(2) - 1 = 2nd term 4(3) - 1 = 3d term So, 4-1= 3 1 st term 8 -1 = 7 2 nd term 12 -1 = 11 3 rd term So the three terms are 3 , 7 , 11

OpenStudy (anonymous):

@Openstudier, Thank you so much :D